# Channel Capacity of Near-Field Multiuser Communications

### v1

@@ IV-B Capacity Region
$\hat{\mathsf{R}}_{1}^{2\rightarrow 1}=\log_{2}\left(1+p_{1}\sigma_{1}^{-2}G_{1}\right),$ (47a)
$\hat{\mathsf{R}}_{2}^{2\rightarrow 1}\!=\log_{2}\!\Big{(}1\!+\frac{p_{2}\sigma_{2}^{-2}G_{2}\!+\!p_{1}p_{2}\sigma_{1}^{-2}\sigma_{2}^{-2}G_{1}G_{2}\bar{\rho}}{1+p_{1}\sigma_{1}^{-2}G_{1}}\Big{)}.$ (47b)
For any given power allocation scheme $\{p_{k}\}_{k=1}^{2}$ with $p_{1}+p_{2}=P$, we can obtain the closed-form expressions for the corner points of the dual MAC regions. The BC capacity region can be then characterized by the convex hull of these corner points.
@@
When $M_{x},M_{z}\rightarrow\infty$, the corner points of the dual MAC capacity region with a given power allocation $\{p_{k}\}_{k=1}^{2}$ satisfy
@@
$~{}~{}~{}\approx\left(\log_{2}(1+{p_{1}\xi}/{(2\sigma_{1}^{2})}),\log_{2}(1+{p_{2}\xi}/{(2\sigma_{2}^{2})})\right).$ (48)
@@
IV-C Comparison with the FF Capacity
The FF BC capacity is given as follows.
@@ IV-C Comparison with the FF Capacity
For the FF case, the asymptotic BC capacity for $M\rightarrow\infty$ is given in (49), shown at the bottom of the next page, which yields $\mathsf{C}_{\mathsf{BC}}^{\mathsf{F}}\simeq\mathcal{O}(\log M)$.
@@
According to the results of (50) below, we obtain the following remarks.
@@
In contrast to the NF case, the FF BC capacity with co-directional UTs is smaller than that when the UTs are in different directions.
The capacity region of the BC under the FF model, when considering an infinitely large array, is influenced by the relative directions of the UTs.
Corollary 10.
In the FF scenario, when $M\rightarrow\infty$, if the UTs are located in the same direction, the corner points of the dual MAC capacity region satisfy $(\hat{\mathsf{R}}_{1}^{1\rightarrow 2},\hat{\mathsf{R}}_{2}^{1\rightarrow 2})\simeq(\log_{2}(1\!+\frac{p_{1}r_{2}^{2}\sigma_{2}^{2}}{p_{2}r_{1}^{2}\sigma_{1}^{2}}),\mathcal{O}(\log M))$ and $(\hat{\mathsf{R}}_{1}^{2\rightarrow 1},\hat{\mathsf{R}}_{2}^{2\rightarrow 1})\simeq(\mathcal{O}(\log M),\log_{2}(1+\frac{p_{2}r_{1}^{2}\sigma_{1}^{2}}{p_{1}r_{2}^{2}\sigma_{2}^{2}}))$. If the UTs are positioned in different directions, we have $(\hat{\mathsf{R}}_{1}^{1\rightarrow 2},\hat{\mathsf{R}}_{2}^{1\rightarrow 2})\simeq(\hat{\mathsf{R}}_{1}^{2\rightarrow 1},\hat{\mathsf{R}}_{2}^{2\rightarrow 1})\simeq(\mathcal{O}(\log M),\mathcal{O}(\log M))$.
Proof:
Similar to the proof of Corollary 4. ∎
Remark 14.
The results in Corollary 10 suggest that the BC capacity region under the FF model can extend unboundedly with the number of the array elements, which is impractical.
Remark 15.
The FF BC capacity region of cases when the UTs are in different directions is more extensive than that of cases when the UTs are in a same direction.
$\mathsf{C}_{\mathsf{BC}}^{\mathsf{F}}\simeq\begin{cases}\log_{2}\!\left(\frac{MPA}{4\pi}\max_{k=1,2}\frac{\Psi_{k}}{r_{k}^{2}\sigma_{k}^{2}}\right)\triangleq\mathsf{F}_{\mathsf{BC}}^{\mathsf{s}}&\left(\theta_{1},\phi_{1}\right)=\left(\theta_{2},\phi_{2}\right)\\ \log_{2}\left(\frac{\left(MPA\left(4\pi\right)^{-1}+\Psi_{1}^{-1}r_{1}^{2}\sigma_{1}^{2}+\Psi_{2}^{-1}r_{2}^{2}\sigma_{2}^{2}\right)^{2}}{4\Psi_{1}^{-1}\Psi_{2}^{-1}r_{1}^{2}r_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2}}-1\right)\triangleq\mathsf{F}_{\mathsf{BC}}^{\mathsf{d}}&\left(\theta_{1},\phi_{1}\right)\neq\left(\theta_{2},\phi_{2}\right)\\ \end{cases}$ (49)
@@ IV-D Extension to Cases of $K>2$
${\mathsf{C}}_{\mathsf{BC}}=\max_{\{{\bm{\Sigma}}_{\varepsilon(k)}\}_{k=1}^{K}:{\bm{\Sigma}}_{\varepsilon(k)}\succeq{\mathbf{0}},\sum_{k=1}^{K}{\mathsf{tr}}({\bm{\Sigma}}_{\varepsilon(k)})\leq P}{\mathsf{R}}_{\varepsilon(k)}$
$=\max_{0\leq p_{k},\sum_{k=1}^{K}p_{k}=P}\log_{2}\det\left({\mathbf{I}}\!+\sum\nolimits_{k=1}^{K}\frac{p_{k}}{\sigma_{k}^{2}}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\right).$ (51)
This is a convex problem which can be optimally solved via the sum power iterative water-filling method [23]. It is easily shown that the results can be expressed in terms of a function of the channel gain and CCF. Furthermore, given a power allocation scheme $\{p_{k}\}_{k=1}^{K}$ with $\sum_{k=1}^{K}p_{k}=P$, we can obtain the capacity region of a dual MAC. The capacity region of the BC is the convex hull of all these dual MAC capacity regions, which means that the boundary of the BC capacity region is formed by the corner points of the dual MAC capacity regions.
@@ V-A NF Multicast Capacity
Although the problem is convex, no closed-form or water filling-based solution is know to exist. However, for practical purposes, the covariance matrix $\bm{\Sigma}$ is often constrained. When the transmit beamforming is employed, we have $\bm{\Sigma}=P\mathbf{w}\mathbf{w}^{\mathsf{H}}$, constraining $\bm{\Sigma}$ to be unit rank [12, 24]. In this case, the multicast capacity in (53) can be reformulated as follows:
@@
${\mathbf{w}}^{\star}=\begin{cases}{G_{1}}^{-\frac{1}{2}}{{\mathbf{h}}_{1}}&{\frac{G_{1}}{\sigma_{1}^{2}}\leq\rho_{{}_{\mathsf{N}}}\frac{G_{2}}{\sigma_{2}^{2}}}\\ {G_{2}}^{-\frac{1}{2}}{{\mathbf{h}}_{2}}&{\frac{G_{2}}{\sigma_{2}^{2}}\leq\rho_{{}_{\mathsf{N}}}\frac{G_{1}}{\sigma_{1}^{2}}}\\ \frac{\mu_{1}}{\sigma_{1}\sqrt{\eta}}{\mathbf{h}}_{1}+\frac{\mu_{2}}{\sigma_{2}\sqrt{\eta}}{\mathbf{h}}_{2}{\rm{e}}^{-{\rm{j}}\angle({\mathbf{h}}_{1}^{\mathsf{H}}{\mathbf{h}}_{2})}&{\text{else}}\end{cases},$ (56)
where $\mu_{1}=\frac{\sigma_{1}^{2}G_{2}-\sigma_{1}\sigma_{2}\sqrt{G_{1}G_{2}\rho_{{}_{\mathsf{N}}}}}{\chi}$, $\mu_{2}=\frac{\sigma_{2}^{2}G_{1}-\sigma_{1}\sigma_{2}\sqrt{G_{1}G_{2}\rho_{{}_{\mathsf{N}}}}}{\chi}$, $\eta=\frac{G_{1}G_{2}\bar{\rho}}{\chi}$, and $\chi=\sigma_{2}^{2}G_{1}+\sigma_{1}^{2}G_{2}-2\sigma_{1}\sigma_{2}\sqrt{G_{1}G_{2}\rho_{{}_{\mathsf{N}}}}$. Accordingly, the multicast capacity is given by
${\mathsf{C}}_{\mathsf{MC}}=\begin{cases}\log_{2}(1\!+\!P\sigma_{1}^{-2}G_{1})&{{\sigma_{1}^{-2}}{G_{1}}\leq\rho_{{}_{\mathsf{N}}}{\sigma_{2}^{-2}}{G_{2}}}\\ \log_{2}(1\!+\!P\sigma_{2}^{-2}G_{2})&{{\sigma_{2}^{-2}}{G_{2}}\leq\rho_{{}_{\mathsf{N}}}{\sigma_{1}^{-2}}{G_{1}}}\\ \log_{2}\left(1+P\eta\right)&{\text{else}}\end{cases}.$ (57)
@@
Please refer to Appendix -F for more details. ∎
@@
$\mathsf{C}_{\mathsf{MC}}^{\mathsf{F}}\simeq\begin{cases}\log_{2}\left(\frac{MPA}{4\pi}\min_{k=1,2}\frac{\Psi_{k}}{r_{k}^{2}\sigma_{k}^{2}}\right)\triangleq\mathsf{F}_{\mathsf{MC}}^{\mathsf{s}}&\left(\theta_{1},\phi_{1}\right)=\left(\theta_{2},\phi_{2}\right)\\ \log_{2}\left(\frac{MPA}{4\pi\left(\Psi_{1}^{-1}r_{1}^{2}\sigma_{1}^{2}+\Psi_{2}^{-1}r_{2}^{2}\sigma_{2}^{2}\right)}\right)\triangleq\mathsf{F}_{\mathsf{MC}}^{\mathsf{d}}&\left(\theta_{1},\phi_{1}\right)\neq\left(\theta_{2},\phi_{2}\right)\\ \end{cases}.$ (61)

### v2

@@ IV-A2 Capacity Region
$\hat{\mathsf{R}}_{1}^{1\rightarrow 2}=\log_{2}\Big{(}1+\frac{\tilde{\gamma}_{2}\mathsf{g}_{2}^{\mathsf{n}}+\tilde{\gamma}_{1}\mathsf{g}_{1}^{\mathsf{n}}\tilde{\gamma}_{2}\mathsf{g}_{2}^{\mathsf{n}}(1-\rho_{\mathsf{n}})}{1+\tilde{\gamma}_{2}\mathsf{g}_{2}^{\mathsf{n}}}\Big{)},$ (61a)
$\hat{\mathsf{R}}_{2}^{1\rightarrow 2}=\log_{2}\left(1+\tilde{\gamma}_{2}\mathsf{g}_{2}^{\mathsf{n}}\right).$ (61b)
For any given power allocation scheme $\{p_{k}\}_{k=1}^{2}$ with $p_{1}+p_{2}=P$, we can obtain the closed-form expressions for the corner points of the dual MAC regions. The NF BC capacity region can be then characterized by the convex hull of these corner points.
@@
When $M_{x},M_{z}\rightarrow\infty$, the corner points of the NF dual MAC capacity region with a given power allocation $\{p_{k}\}_{k=1}^{2}$ satisfy
@@
$~{}~{}~{}\approx\left(\log_{2}(1+{\tilde{\gamma}_{1}\xi}/{2}),\log_{2}(1+{\tilde{\gamma}_{2}\xi}/{2})\right).$ (62)
@@
IV-B Comparison with the FF Capacity
Next, we discuss the FF BC capacity for comparison.
@@ IV-B Comparison with the FF Capacity
For the FF case, the asymptotic BC capacity for $M\rightarrow\infty$ is given in (63), shown at the bottom of this page, which yields $\mathsf{C}_{\mathsf{BC}}\simeq\mathcal{O}(\log M)$.
@@
Further, according to the results of (64), we find that, in contrast to the NF case, the FF BC capacity with co-directional UTs is smaller than that when the UTs are in different directions.
The capacity region of the BC in the FF scenario, when considering an infinitely large array, is influenced by the relative directions of the UTs. Specifically, when $M\rightarrow\infty$, if the UTs are located in the same direction, the corner points of the dual MAC capacity region satisfy
$(\hat{\mathsf{R}}_{1}^{1\rightarrow 2},\hat{\mathsf{R}}_{2}^{1\rightarrow 2})$ $\simeq(\log_{2}(1+{\tilde{\gamma}_{1}r_{2}^{2}}/{(\tilde{\gamma}_{2}r_{1}^{2})}),\mathcal{O}(\log M)),$ (65a)
$(\hat{\mathsf{R}}_{1}^{2\rightarrow 1},\hat{\mathsf{R}}_{2}^{2\rightarrow 1})$ $\simeq(\mathcal{O}(\log M),\log_{2}(1\!+{\tilde{\gamma}_{2}r_{1}^{2}}/{(\tilde{\gamma}_{1}r_{2}^{2})})).$ (65b)
If the UTs are positioned in different directions, we have
$\!\!(\hat{\mathsf{R}}_{1}^{1\rightarrow 2}\!,\hat{\mathsf{R}}_{2}^{1\rightarrow 2})\!\approx\!(\hat{\mathsf{R}}_{1}^{2\rightarrow 1}\!,\hat{\mathsf{R}}_{2}^{2\rightarrow 1})\!\simeq\!(\mathcal{O}(\log M),\mathcal{O}(\log M)).\!\!$ (66)
The above results suggest that the BC capacity region under the FF model can extend unboundedly with the number of the array elements, which is impractical. Additionally, the FF BC capacity region of cases when the UTs are in different directions is more extensive than that of cases when the UTs are in a same direction.
IV-C Comparison with Linear Precoders
[…]
$\hat{\mathsf{R}}_{k}=\log_{2}\bigg{(}1+\frac{\lvert{\mathbf{w}}_{k}^{\mathsf{H}}{\mathbf{h}}_{k}\rvert^{2}\hat{p}_{k}}{\sigma_{k}^{2}+\sum_{k^{\prime}\neq k}\lvert{\mathbf{w}}_{k^{\prime}}^{\mathsf{H}}{\mathbf{h}}_{k}\rvert^{2}\hat{p}_{k^{\prime}}}\bigg{)},$ (67)
where $\hat{p}_{k}$ is the transmit power. For the sake of brevity, we continue to consider the two-user case with $K=2$.
IV-C1 Maximal-Ratio Transmission Precoding
We begin with the maximal-ratio transmission (MRT) precoder, which is designed to maximize the channel gain for each UT. This yields ${\mathbf{w}}_{k}=\frac{{\mathbf{h}}_{k}}{\lVert{\mathbf{h}}_{k}\rVert}$ for $k\in{\mathcal{K}}$. Substituting ${\mathbf{w}}_{k}=\frac{{\mathbf{h}}_{k}}{\lVert{\mathbf{h}}_{k}\rVert}$ into (67) and performing some basic mathematical manipulations, we express the sum-rate as follows:
$\hat{\mathsf{R}}_{1}+\hat{\mathsf{R}}_{2}=\sum_{k=1}^{2}\log_{2}\Big{(}1+{\hat{\gamma}}_{k}{\mathsf{g}}_{k}\Big{(}1-\frac{\hat{\gamma}_{k^{\prime}}{\mathsf{g}}_{k}\rho}{1+\hat{\gamma}_{k^{\prime}}{\mathsf{g}}_{k}\rho}\Big{)}\Big{)}\triangleq\hat{\mathsf{R}}_{{\mathsf{mrt}}},$ (68)
where ${\hat{\gamma}}_{k}=\frac{\hat{p}_{k}}{\sigma_{k}^{2}}$ denotes the SNR.
IV-C2 Zero-Forcing Precoding
Next, we consider ZF precoding, which aims to nullify the IUI. In this case, ${\mathbf{w}}_{k}^{\mathsf{H}}{\mathbf{h}}_{k^{\prime}}=0$ for $k^{\prime}\neq k$. Following the steps to obtain (37), we write the ZF precoder for UT $k$ as follows:
${\mathbf{w}}_{k}=\frac{(\mathbf{I}_{M}-\mathbf{h}_{k^{\prime}}\mathbf{h}_{k^{\prime}}^{\mathsf{H}}/\|\mathbf{h}_{k^{\prime}}\|^{2})\mathbf{h}_{k}}{\lVert(\mathbf{I}_{M}-\mathbf{h}_{k^{\prime}}\mathbf{h}_{k^{\prime}}^{\mathsf{H}}/\|\mathbf{h}_{k^{\prime}}\|^{2})\mathbf{h}_{k}\rVert}.$ (69)
Substituting (69) into (67) and performing some simple manipulations, we obtain
$\hat{\mathsf{R}}_{1}+\hat{\mathsf{R}}_{2}=\sum\nolimits_{k=1}^{2}\log_{2}(1+\hat{\gamma}_{k}\mathsf{g}_{k}(1-\rho))\triangleq{\hat{\mathsf{R}}}_{{\mathsf{zf}}}.$ (70)
IV-C3 Summary and Discussion
Through the derivations above, we observe that the sum-rate achieved by the considered linear precoding schemes can be expressed as a function of the channel gains and the CCF, i.e.,
$\hat{\mathsf{R}}_{\ell}=\sum\nolimits_{k=1}^{2}\log_{2}(1+\hat{\gamma}_{k}\mathsf{g}_{k}(1-\hat{f}_{\ell}(\hat{\gamma}_{k^{\prime}}\mathsf{g}_{k},\rho))),$ (71)
where $\ell\in\{{\mathsf{mrt}},{\mathsf{zf}}\}$, $\hat{f}_{{\mathsf{mrt}}}(x,z)\triangleq\frac{xz}{1+xz}$, and $\hat{f}_{{\mathsf{zf}}}(x,z)\triangleq z$. By substituting (15) and (16) into (71), we obtain the NF sum-rate expressions under the linear precoders, which can be written as follows:
$\hat{\mathsf{R}}_{\ell}=\sum\nolimits_{k=1}^{2}\log_{2}(1+\hat{\gamma}_{k}\mathsf{g}_{k}^{\mathsf{n}}(1-\hat{f}_{\ell}(\hat{\gamma}_{k^{\prime}}\mathsf{g}_{k}^{\mathsf{n}},\rho_{\mathsf{n}}))).$ (72)
As previously mentioned, we have $\lim_{M_{x},M_{z}\rightarrow\infty}\rho\approx 0$ under the NF channel model. This means that
$\lim_{M_{x},M_{z}\rightarrow\infty}\hat{\mathsf{R}}_{\ell}\approx\sum\nolimits_{k=1}^{2}{\log_{2}\left(1+\hat{\gamma}_{k}\mathsf{g}_{k}^{\mathsf{n}}\right)}$ (73)
holds for $\ell\in\{{\mathsf{mrt}},{\mathsf{zf}}\}$.
@@ IV-C3 Summary and Discussion
The above discussion implies that in NF BC, when the BS antenna number is sufficiently large, the considered linear precoding schemes can approach the upper bound of the sum-rate capacity, akin to the capacity-achieving DPC.
(a) $(\theta_{2},\phi_{2})=(\frac{\pi}{3},\frac{2\pi}{3})$.
(b) $(\theta_{2},\phi_{2})=(\frac{\pi}{3}\!+\!\frac{\pi}{60},\frac{2\pi}{3}\!+\!\frac{\pi}{60})$.
Figure 5: $\Gamma_{{\mathsf{dl}},\ell}$ versus $M$ for different precoders with $(\theta_{1},\phi_{1})=(\frac{\pi}{3},\frac{2\pi}{3})$ and the other parameter settings given in Section VI.
This conclusion in Remark 9 holds when $M_{x},M_{z}\rightarrow\infty$. However, in practical NFC systems with large yet finite values of $M_{x},M_{z}$, there will be a gap between the achievable sum-rate and its bound $\sum_{k=1}^{2}{\log_{2}\left(1+\hat{\gamma}_{k}\mathsf{g}_{k}^{\mathsf{n}}\right)}$, which is given by
$\Gamma_{{\mathsf{dl}},\ell}\triangleq\frac{\sum\nolimits_{k=1}^{2}\log_{2}(1+\hat{\gamma}_{k}\mathsf{g}_{k}^{\mathsf{n}}(1-\hat{f}_{\ell}(\hat{\gamma}_{k^{\prime}}\mathsf{g}_{k}^{\mathsf{n}},\rho_{\mathsf{n}})))}{\sum_{k=1}^{2}{\log_{2}\left(1+\hat{\gamma}_{k}\mathsf{g}_{k}^{\mathsf{n}}\right)}}.$ (74)
Figure 5 plots $\Gamma_{{\mathsf{dl}},\ell}$ versus the BS antenna number $M$ under equal power allocation, i.e., $\hat{p}_{k}=\frac{P}{K}$ for $k=1,2$. For comparison, the ratio between the rate achieved by DPC precoder and the upper bound is also presented. As shown, with the increasing number of antennas, the considered precoding schemes achieve a sum-rate close to their upper bound. Among the linear schemes, the MRT precoding scheme achieves superior sum-rate performance for smaller $M$, while NF precoder is more effective with large-scale array. Additionally, we can observed that linear precoders exhibit comparable performance to the capacity-achieving DPC precoders for larger $M$ in the NF scenario.
@@ IV-D Extension to Cases of $K>2$
${\mathsf{C}}_{\mathsf{BC}}=\max_{0\leq p_{k},\sum_{k=1}^{K}p_{k}=P}\log_{2}\det\left({\mathbf{I}}+\sum\nolimits_{k=1}^{K}\frac{p_{k}}{\sigma_{k}^{2}}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\right)$
[…]
Note that calculating the BC capacity ${\mathsf{C}}_{\mathsf{BC}}$ involves solving the convex optimization problem formulated on the right-hand side of (75), which can be efficiently addressed using the sum-power iterative water-filling method, as introduced in [32]. It is clearly shown that the results can be expressed in terms of the channel gains and CCFs. Furthermore, given a power allocation scheme $\{p_{k}\}_{k=1}^{K}$ with $\sum_{k=1}^{K}p_{k}=P$, we can obtain the capacity region of a dual MAC. The capacity region of the BC is the convex hull of all these dual MAC capacity regions, which means that the boundary of the BC capacity region is formed by the corner points of the dual MAC capacity regions.
@@ V-A NF Multicast Capacity
This suggests that the multicast capacity can be obtained by solving a convex problem, as shown on the right-hand side of (77). However, no closed-form or water-filling-based solution is known to exist for this problem.
For practical purposes, the covariance matrix $\bm{\Sigma}$ is often constrained. When transmit beamforming is employed, it holds that $\bm{\Sigma}=P\mathbf{w}\mathbf{w}^{\mathsf{H}}$, which constrains $\bm{\Sigma}$ to be of unit rank [18, 33]. In this case, the multicast capacity in (77) can be reformulated as follows:
@@
${\mathbf{w}}^{\star}=\begin{cases}{\mathsf{g}_{1}^{\mathsf{n}}}^{-\frac{1}{2}}{{\mathbf{h}}_{1}}&{\frac{\mathsf{g}_{1}^{\mathsf{n}}}{\sigma_{1}^{2}}\leq\rho_{\mathsf{n}}\frac{\mathsf{g}_{2}^{\mathsf{n}}}{\sigma_{2}^{2}}}\\ {\mathsf{g}_{2}^{\mathsf{n}}}^{-\frac{1}{2}}{{\mathbf{h}}_{2}}&{\frac{\mathsf{g}_{2}^{\mathsf{n}}}{\sigma_{2}^{2}}\leq\rho_{\mathsf{n}}\frac{\mathsf{g}_{1}^{\mathsf{n}}}{\sigma_{1}^{2}}}\\ \frac{\mu_{1}}{\sigma_{1}\sqrt{\eta}}{\mathbf{h}}_{1}+\frac{\mu_{2}}{\sigma_{2}\sqrt{\eta}}{\mathbf{h}}_{2}{\rm{e}}^{-{\rm{j}}\angle({\mathbf{h}}_{1}^{\mathsf{H}}{\mathbf{h}}_{2})}&{\text{else}}\end{cases},$ (80)
where $\mu_{1}=\frac{\sigma_{1}^{2}\mathsf{g}_{2}^{\mathsf{n}}-\sigma_{1}\sigma_{2}\sqrt{\mathsf{g}_{1}^{\mathsf{n}}\mathsf{g}_{2}^{\mathsf{n}}\rho_{\mathsf{n}}}}{\chi}$, $\mu_{2}=\frac{\sigma_{2}^{2}\mathsf{g}_{1}^{\mathsf{n}}-\sigma_{1}\sigma_{2}\sqrt{\mathsf{g}_{1}^{\mathsf{n}}\mathsf{g}_{2}^{\mathsf{n}}\rho_{\mathsf{n}}}}{\chi}$, $\eta=\frac{\mathsf{g}_{1}^{\mathsf{n}}\mathsf{g}_{2}^{\mathsf{n}}(1-\rho_{\mathsf{n}})}{\chi}$, and $\chi=\sigma_{2}^{2}\mathsf{g}_{1}^{\mathsf{n}}+\sigma_{1}^{2}\mathsf{g}_{2}^{\mathsf{n}}-2\sigma_{1}\sigma_{2}\sqrt{\mathsf{g}_{1}^{\mathsf{n}}\mathsf{g}_{2}^{\mathsf{n}}\rho_{\mathsf{n}}}$. Accordingly, the multicast capacity is given by
${\mathsf{C}}_{\mathsf{MC}}=\begin{cases}\log_{2}(1\!+\!P\sigma_{1}^{-2}\mathsf{g}_{1}^{\mathsf{n}})&{{\sigma_{1}^{-2}}{\mathsf{g}_{1}^{\mathsf{n}}}\leq\rho_{\mathsf{n}}{\sigma_{2}^{-2}}{\mathsf{g}_{2}^{\mathsf{n}}}}\\ \log_{2}(1\!+\!P\sigma_{2}^{-2}\mathsf{g}_{2}^{\mathsf{n}})&{{\sigma_{2}^{-2}}{\mathsf{g}_{2}^{\mathsf{n}}}\leq\rho_{\mathsf{n}}{\sigma_{1}^{-2}}{\mathsf{g}_{1}^{\mathsf{n}}}}\\ \log_{2}\left(1+P\eta\right)&{\text{else}}\end{cases}.$ (81)
@@
Please refer to Appendix -D for more details. ∎
@@
Figure 7: MAC sum-rate capacity versus $M$.
Figure 8: BC sum-rate capacity versus $M$.